# Survey of Computerized Adaptive Testing: A Machine Learning Perspective

### 05 Apr 2024 v2

@@ Abstract.
Computerized Adaptive Testing (CAT) provides an efficient and tailored method for assessing the proficiency of examinees, by dynamically adjusting test questions based on their performance. Widely adopted across diverse fields like education, healthcare, sports, and sociology, CAT has revolutionized testing practices. While traditional methods rely on psychometrics and statistics, the increasing complexity of large-scale testing has spurred the integration of machine learning techniques. This paper aims to provide a machine learning-focused survey on CAT, presenting a fresh perspective on this adaptive testing method. By examining the test question selection algorithm at the heart of CAT’s adaptivity, we shed light on its functionality. Furthermore, we delve into cognitive diagnosis models, question bank construction, and test control within CAT, exploring how machine learning can optimize these components. Through an analysis of current methods, strengths, limitations, and challenges, we strive to develop robust, fair, and efficient CAT systems. By bridging psychometric-driven CAT research with machine learning, this survey advocates for a more inclusive and interdisciplinary approach to the future of adaptive testing.
Computerized Adaptive Testing, Machine Learning, Cognitive Diagnosis, Intelligent Measurement, Question Selection
ccs: Applied computing Learning management systemsccs: Applied computing Computer-managed instruction
1. Introduction
Computerized Adaptive Testing (CAT), originating from psychometrics, allows for efficient assessment through personalized question selection (Chen et al., 2015a; Vie et al., 2017). It has been widely applied in high-stakes testing scenarios, including the Duolingo Test, GRE, TOEFL, and GMAT. The term “computerized” underscores its origins, reflecting a combination of computer science and intelligent measurement. As shown in Figure 1, in contrast to the one-for-all approach of traditional paper-and-pencil tests, CAT tailors the selection of questions to each examinee’s level of proficiency, thereby maximizing the accuracy of the assessment while minimizing the test length. In essence, CAT strives to answer a pivotal question about accuracy and efficiency: How to accurately estimate an examinee’s true proficiency by providing the minimal number of questions?
Figure 1. If (a) traditional paper-and-pencil test is “one-for-all”, then (b) CAT is “one-for-each”. Each examinee gets a personalized test that adapts to his/her proficiency level and knowledge, ensuring each question accurately assesses and challenges the examinee
@@ 1. Introduction
CAT is a dynamic and interactive process between an examinee and a testing system. The CAT system includes four main components that take turns: At each test step, the Cognitive Diagnosis Model, as the user model, first uses the examinee’s previous responses to estimate their current proficiency, based on cognitive science or psychometrics (Ackerman et al., 2003). Then, the Selection Algorithm picks the next question from the Question Bank according to certain criteria (Lord, 2012; Chang and Ying, 1996; Bi et al., 2020). Most traditional criteria are statistical informativeness metrics, e.g., selecting the question whose difficulty matches the examinee’s current proficiency estimate, meaning the examinee has roughly a 50% chance of getting it right. The above process repeats until a predefined stopping rule is met. Throughout the assessment, Test Control governs various factors such as content balance, fairness, and robustness of the testing process. At the conclusion of the CAT, the final proficiency estimate—or the diagnostic report—is provided to the examinee as the outcome of this assessment.
@@
Obviously, CAT represents a complex fusion of machine intelligence and assessment techniques. It needs to manage large question banks, adapt to varying examinee proficiencies, and real-time decision-making. Moreover, practical CAT also involves ensuring reliability, fairness, search efficiency, etc. These challenges make CAT a multifaceted decision-making problem that requires sophisticated solutions. Particularly, with the emergence of large-scale and diverse online testing platforms, traditional statistical methods often struggle to optimally select questions while addressing these multifarious factors. Machine learning (Jordan and Mitchell, 2015), with its ability to learn from data and make precise predictions, offers a viable solution to enhance test efficiency and accuracy within this volatile environment. Previous CAT surveys (Chang, 2015; Vie et al., 2017; Cheng, 2008; Mujtaba and Mahapatra, 2020) have largely revolved around the domains of statistics and psychometrics. Given CAT’s interdisciplinary nature, this paper seeks to understand and review methodologies from a machine-learning perspective, which is more approachable to a wider readership and sheds light on building strong CAT systems.
@@
In the world of machine learning, CAT can be viewed as a parameter estimation problem with a focus on data efficiency (Mirzasoleiman et al., 2020; Zhuang et al., 2023): the goal is to sequentially select the smallest number of data samples (questions) from a pool of unlabeled data (the question bank) so that, after their labels (correct/wrong responses) are observed, the model’s latent parameters (examinee’s true proficiency $\theta_{0}$) can be estimated accurately. In recent years, there has been a growing interest in applying machine learning techniques to investigate the four components in CAT. For example, deep learning techniques diagnose examinee’s proficiency (Wang et al., 2020c) and automate question bank construction (Qiu et al., 2019); data-driven approaches optimize selection algorithms by learning from large-scale response data (Zhuang et al., 2022b; Li, 2020; Ghosh and Lan, 2021); machine learning enhances CAT’s reliability by addressing various factors (e.g., robustness, search efficiency) (Hong et al., 2023). Despite these efforts, a comprehensive survey that captures the breadth of CAT solutions from a machine-learning perspective is still lacking. Furthermore, the ongoing evolution of machine learning presents new aspects for CAT.
This paper will help identify the strengths and limitations of machine learning approaches in CAT, highlight the challenges, and provide guidance for future research directions. Given the fact that the selection algorithm is the core component for achieving the adaptivity, this survey mainly focuses on its advancements. The contributions of this paper are as follows:
@@
  • To our knowledge, this represents the first attempt to comprehensively review CAT solutions through the lens of machine learning. By exploring the existing work in Cognitive Diagnosis Model, Selection Algorithms, Question Bank Construction, and Test Control, the paper offers a unified framework and encompasses the entire life cycle of the CAT system.
  • We summarize existing works and draw conclusions on the success and failure attempts of machine learning. We also discuss the crucial factors in building a reliable CAT, e.g., content balance, fairness, robustness, and search efficiency, often overlooked in prior reviews.
@@
The paper is organized as follows. In Section 2, we introduce the history and the contemporary trend of CAT. Then in Section 3, we provide the formulation of CAT’s task. After that, Section 4, 5, and 6 respectively review the existing methods for the cognitive diagnosis model, selection algorithms, and question bank construction. In Section 7, we summarize the key factors in the application of CAT. Section 8 discusses how to evaluate the CAT. We discuss the future directions in Section 9 and Section 10 concludes the paper.
2. Background of CAT
Evolution of CAT
The evolution of CAT is a fascinating journey through time, marked by significant milestones that have shaped the landscape of educational measurement.
  • The Rise of CAT. The story begins in the early 20th century with Alfred Binet, a psychologist who developed the first form of adaptive testing in 1905 as an intelligence test (Wainer et al., 2000). In the 1950s, computers began transforming adaptive testing into what we now call CAT. Groundbreaking work throughout the 1970s and 1980s solidified the foundation of CAT, with subsequent refinements enhancing precision and efficiency (Sands et al., 1997). A pivotal advancement was the integration of the Rasch model from Item Response Theory (IRT) (Roskam and Jansen, 1984), which allowed CAT to adaptively match question difficulty to a test-taker’s estimated proficiency level, thereby improving the accuracy of proficiency assessments.
  • Widespread Adoption. In the 1990s, the advent of the internet made CAT more accessible, leading to its adoption by major tests like the GRE and TOEFL. While these tests have evolved from the classic CAT paradigm, adaptive testing principles continue to inform their current testing models. Various statistical methods—Bayesian, maximum likelihood, or maximum information methods–are employed to optimize the testing experience. During this time, CAT emerged as a dominant topic in educational measurement forums. For example, at the 1999 annual meeting of the National Council on Measurement in Education, a quarter of the papers presented were devoted to CAT (Meijer and Nering, 1999).
@@ 2. Background of CAT
  • CAT Meets Machine Learning. Current CAT research spans a wide range of topics, including the development of question banks, question selection, proficiency estimation, and various issues related to test security and reliability. They are critical to ensuring that CAT remains a reliable, valid, and fair assessment. Machine Learning (ML), particularly through its branches such as deep learning and reinforcement learning, is revolutionizing CAT by enabling sophisticated analysis of large datasets, detailed behavior modeling, and flexible adaptation to diverse testing environments (Zhuang et al., 2022b). Despite the ML in CAT is still in its early stages, its potential is evident, with deep learning, natural language processing, and reinforcement learning showing significant promise in enhancing question characterization, scoring processes, and adaptive modeling (Mujtaba and Mahapatra, 2020). This survey aims to provide an overview and understanding of traditional statistical-based and recent ML-based CAT.
Application of CAT
CAT has transcended the boundaries of educational assessment, finding its place in diverse fields such as healthcare, sociology, and sports (Yurtcu and GÜZELLER, 2021).
  • Education. CAT, with its roots in the U.S. Department of Defense’s ASVAB test in the 1980s, has transformed education assessments. It personalizes tests for each individual, benefiting fields such as nursing licensing exams, graduate school exams (GRE and GMAT), and mathematics courses (Verschoor and Straetmans, 2010). There’s also a newer CAT variant called Multi-Stage Testing (MST). While traditional CAT adapts question by question, MST adapts in stages, with a set of questions being provided at each stage (Luecht and Nungester, 1998; Wainer and Kiely, 1987). MST is gaining popularity in exams like the GRE, law school exams (LSAC), and accounting exams (CPA).
  • Healthcare. In healthcare, particularly in mental health, there’s a crucial need for quick and accurate assessments. CAT meets this demand by offering quick evaluations akin to those conducted by skilled clinicians, thereby enhancing the overall efficiency of medical care. CAT for mental health assessments creates scales, including those for depression, anxiety, and suicide risk (Gibbons et al., 2018; Gibbons and deGruy, 2019; Gibbons et al., 2017, 2014), have shown promising results compared to traditional methods (Bagby et al., 2004).
  • Sociology. In the domain of sociology, CAT can be used in public opinion surveys. Montgomery and Cutler (2013) introduce an algorithm that adaptively selects survey questions. It not only enhances the accuracy of the survey results but also significantly reduces the number of questions needed, thereby streamlining the process and cutting down on associated costs.
  • Sports. The evaluation of specific skills is crucial for the development of athletes. Notably, Ando et al. (2018) innovatively apply CAT to the realm of football. They create a standardized assessment called the TSS-CAT, which utilizes video recordings of football tactics as test questions to gauge a player’s tactical proficiency. This groundbreaking assessment tool offers a tailored and efficient way for players, coaches, and managers to monitor and evaluate the progression of individual tactical skills throughout the training process.
3. Overview
An important assumption (Chang, 2015) of CAT is that examinee’s true proficiency level $\theta_{0}\in\mathbb{R}^{d}$ is constant throughout the test, where $d$ refers to the proficiency’s dimension; for example, $\theta$ could denote an overall ability level ($d=1$) or represent the degree of mastery across $d$ knowledge concepts. In CAT, the aim is to estimate examinees’ true proficiency levels (i.e., $\theta_{0}$) by having them answer questions. The ultimate objectives of CAT are twofold: (1) to use the responses to estimate an examinee’s proficiency $\theta$ such that it closely approximates the true proficiency $\theta_{0}$ by the end of the test, and (2) to select the most valuable and fitting questions for each examinee, thereby reducing test length.
@@ 3. Overview
Figure 2. The workflow of CAT: At step $t$, the selection algorithm adaptively selects next question $q_{t+1}$ based on examinee’s current proficiency ${\theta}^{t}$ estimated by CDM.
3.1. Task Formalization
@@
As illustrated in Figure 2, at test step $t\in[1,2,...,T]$ in CAT, examinee’s current proficiency estimate $\hat{\theta}^{t}$ is estimated using previous $t$ responses; then leverage $\hat{\theta}^{t}$ to retrieve the next question $q_{t+1}$ from question bank $\mathcal{Q}$ to ask examinee, and receive his/her next response $y_{t+1}$. These interactions form a response sequence $\{(q_{1},y_{1}),(q_{2},y_{2}),...,(q_{T},y_{T})\}$, where $y_{t}=1$ if the response to $q_{t}$ is correct and 0 otherwise. To achieve the goals of CAT, the following two steps are necessary at each test step: Step 1: Proficiency Estimation. The Cognitive Diagnosis Model (CDM), denoted by $f(\cdot)$, acts as a user model, predicting the probability of a correct response by an examinee with proficiency $\theta$, which is denoted as $f(q,\theta)=P(y=1|q,\theta)$. The implementation of CDM often draws upon cognitive science (De La Torre, 2009) or psychometrics(Ackerman et al., 2003). To accurately estimate his/her proficiency at each step, various estimation methods can be used, e.g., Maximum Likelihood Estimation (MLE) or Bayesian Estimation. In applications, the binary cross-entropy loss is frequently utilized: at step $t$, given previous $t$ responses $X_{1:t}=\{(q_{1},y_{1}),(q_{2},y_{2}),...,(q_{t},y_{t})\}$, the corresponding empirical loss is:
(1) ${L}(\theta)=\sum_{(q,y)\in X_{1:t}}l(y,f(q,\theta))=-\sum_{(q,y)\in X_{1:t}}{y\log f(q,\theta)+(1-y)\log(1-f(q,\theta))},$
@@
thus the current estimate of proficiency, $\hat{\theta}^{t}$, is obtained by minimizing the loss function $L(\theta)$: $\hat{\theta}^{t}=\mathop{\arg\min}_{\theta}{{L}(\theta)}$. Step 2: Question Selection. The heart of CAT is an algorithm that picks the next question $q_{t+1}$ from the question bank $\mathcal{Q}$, using examinee’s current proficiency estimate $\hat{\theta}^{t}$ as a guide:
(2) $q_{t+1}=\mathop{\arg\max}_{q\in\mathcal{Q}}\mathcal{V}_{q}(\hat{\theta}^{t}),$
@@
where $\mathcal{V}_{q}(\hat{\theta}^{t})$ is the value of question $q$. For instance, $\mathcal{V}$ might be a measure of how much information the question will provide about the examinee’s proficiency, or it could be the output of a policy $\pi$ specifically designed to determine question selection. After receiving new response $y_{t+1}$, CDM updates and estimates proficiency $\hat{\theta}^{t+1}$. The above process will be repeated for $T$ times, ensuring the final step estimate $\hat{\theta}^{T}$ close to the true $\theta_{0}$, i.e.,
Definition 0 (Definition of CAT).
@@
(3) $\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|.$
However, solving this optimization problem directly is impractical, as the true proficiency $\theta_{0}$ is not observable, and even the examinees may not know their exact proficiency level. Consequently, existing methods are all approximations of this target. For instance, traditional statistical selection methods (Lord, 2012; Chang and Ying, 1996) utilize the asymptotic statistical properties of the MLE to reduce estimation uncertainty, such as choosing questions that closely match the examinee’s current estimated proficiency $\hat{\theta}^{t}$. More recent Subset Selection approaches (Zhuang et al., 2023) try to identify a theoretical approximation of $\theta_{0}$ to serve as a new objective for optimization. For further details, refer to Section 5. Meanwhile, as a practical system, considerations extend beyond this proficiency estimation objective (Definition 1). Factors such as CAT’s question exposure control, robustness, fairness, and search efficiency must be addressed as well. An exhaustive discussion of these factors is presented in Section 7. Evaluation Methods: To validate the accuracy of the estimated proficiency, two primary approaches are employed: 1) using the examinee’s estimated values within the CDM to predict the correctness label $y$ of the responses on his/her reserved response data, often measured by cross-entropy; 2) using simulation to generate true proficiency values $\theta_{0}$, simulating the examinee’s responses to each question. Then the Mean Squared Error (MSE) between the estimates and the simulated true values can be calculated. The details can be found in Section 8.
@@
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=14em,font=,, where level=2text width=15em,font=,, where level=3text width=17em,font=,, [ Computerized Adaptive Testing, ver [ Cognitive Diagnosis Model
     Section 4, fill=red!10, text width=15em [ Latent Trait Model (4.1), fill=red!10, text width=19em [  IRT (Embretson and Reise, 2013; Hambleton et al., 1991; Lord, 2012), MIRT (Reckase, 2006), VI-IRT (Wu et al., 2020)., leaf, text width=22em ] ] [ Diagnostic Classification Model (4.2), fill=red!10, text width=19em [  DINA (De La Torre, 2009; Von Davier, 2014), G-DINA (de la Torre, 2011),
 AHM (Leighton et al., 2004), FuzzyCDF (Liu et al., 2018)., leaf, text width=22em ] ] [ Deep Learning Model (4.3), fill=red!10, text width=19em [  DIRT (Cheng et al., 2019), NeuralCD (Wang et al., 2020c), IRR (Tong et al., 2021),
 RCD (Gao et al., 2021), HierCDF (Li et al., 2022), DeepCDF (Gao et al., 2022),
 BETA-CD (Bi et al., 2023), AGCDM (Pei et al., 2022), SCD (Wang et al., 2023b).,leaf, text width=22em ] ] ] [ Selection Algorithm
   Section 5, fill=blue!10, text width=15em [ Statistical Algorithms (5.1), fill=blue!10, text width=18em [  Classic Information: FIsher Info (Lord, 2012), KL Info (Chang and Ying, 1996; Chang, 2015);
 Additional Information: MPWI (Barrada et al., 2009a), MEI (van der Linden, 1998), thOpt (Barrada et al., 2006);
 CD-CAT: KL algorithm (Cheng, 2009), Shannon entropy (Tatsuoka, 2002);
 Dual-objective CD-CAT (Kang et al., 2017; Zheng et al., 2018).,leaf, text width=27em ] ] [ Active Learning Algorithms (5.2), fill=blue!10, text width=18em [  Uncertainty method (Veldkamp and Verschoor, 2019), Expected Model Change (Bi et al., 2020)., leaf, text width=27em ] ] [ Reinforcement Learning (5.3), fill=blue!10, text width=18em [  SSP (Gilavert and Freire, 2022), POMDP (Nurakhmetov, 2019; Gilavert and Freire, 2022; Chen et al., 2018), DQN (Li et al., 2023),
 GMOCAT (Wang et al., 2023a), NCAT (Zhuang et al., 2022b)., leaf, text width=27em ] ] [ Meta Learning Algorithms (5.4), fill=blue!10, text width=18em [  BOBCAT (Ghosh and Lan, 2021; Feng et al., 2023), DL-CAT (Ma et al., 2023), SACAT (Yu et al., 2023)., leaf, text width=27em ] ] [ Subset Selection Algorithms (5.5), fill=blue!10, text width=18em [  BECAT (Zhuang et al., 2023), SEM (Mujtaba and Mahapatra, 2021)., leaf, text width=27em ] ] ] [ Question Bank Construction
     Section 6, fill=yellow!10, text width=15em [ Question Characteristics Analysis (6.1), fill=yellow!10, text width=20em [  Expert-based (López-Cuadrado et al., 2008; Kozierkiewicz-Hetmańska and Poniatowski, 2014);
 Statistic-based (Chen et al., 2004; Magno, 2009; Chang and Yang, 2009; Sun et al., 2014; Xiong et al., 2022);
 Deep Learning: CNN (Huang et al., 2017), RNN (Qiu et al., 2019; Huang et al., 2019; Lei et al., 2021),
 NLP (Yin et al., 2019; Ning et al., 2023). , leaf, text width=22em ] ] [ Question Bank Development (6.2), fill=yellow!10, text width=20em [  Expert-based (Reckase, 2010; He and Reckase, 2014; Van Der Linden et al., 2000; Way et al., 2005; van der Linden et al., 2006; Ariel et al., 2004; Gulliksen, 2013; Swanson and Stocking, 1993)., leaf, text width=22em ] ] ] [ Test Control
  Section 7, fill=green!10, text width=15em [ Content Control (7.1), fill=green!10, text width=13em [  Sympson-Hetter (Sympson and Hetter, 1985; van der Linden and Veldkamp, 2004, 2007; Barrada et al., 2009b);
 A-Stratified (Chang and Ying, 1999; Chang et al., 2001; Yi and Chang, 2001; Barrada et al., 2006, 2014)., leaf, text width=20em ] ] [ Fairness (7.2), fill=green!10, text width=13em [  Bias in CDMs (Thompson, [n.d.]; Kizilcec and Lee, 2022);
 Bias in Question Bank (Hambleton et al., 1991; Chu and Lai, 2013; Camilli and Shepard, 1994; Mellenbergh, 1989; Roberts, [n.d.]);
 Bias in Selection Algorithms (Lord, 2012);
 Equating (Brigman and Bashaw, 1976; van der Linden, 2000; JAN, [n.d.]; Sawaki, 2001).,leaf, text width=20em ] ] [ Robustness (7.3), fill=green!10, text width=13em [  Robust CAT(Veldkamp and Verschoor, 2019; Zhuang et al., 2022a).,leaf, text width=20em ] ] [ Search Efficiency (7.4), fill=green!10, text width=13em [  PSO (Huang et al., 2009), SECAT (Hong et al., 2023).,leaf, text width=20em ] ] ] ]
@@
Figure 3. Summary of Computerized Adaptive Testing methods in machine learning perspective.
3.2. Categorization

### 09 Mar 2026 v3

@@ Abstract
Computerized Adaptive Testing (CAT) offers an efficient and personalized method for assessing examinee proficiency by dynamically adjusting test questions based on individual performance. Compared to traditional, non-personalized testing methods, CAT requires fewer questions and provides more accurate assessments. As a result, CAT has been widely adopted across various fields, including education, healthcare, sports, sociology, and the evaluation of AI models. While traditional methods rely on psychometrics and statistics, the increasing complexity of large-scale testing has spurred the integration of machine learning techniques. This paper aims to provide a machine learning-focused survey on CAT, presenting a fresh perspective on this adaptive testing paradigm. We delve into measurement models, question selection algorithm, bank construction, and test control within CAT, exploring how machine learning can optimize these components. Through an analysis of current methods, strengths, limitations, and challenges, we strive to develop robust, fair, and efficient CAT systems. By bridging psychometric-driven CAT research with machine learning, this survey advocates for a more inclusive and interdisciplinary approach to the future of adaptive testing.
Index Terms:
Adaptive testing, machine learning, proficiency assessment, AI evaluation, deep learning.
1 Introduction
The assessment of intelligent agents, whether human or AI systems, is essential for ensuring that individuals are well-prepared to meet the demands of their respective roles [112, 134]. For humans, assessment results can determine eligibility for opportunities such as admissions or employment. For AI models, these results can indicate whether a system is suitable for deployment and capable of making real-world decisions. Traditionally, assessments have often used a one-size-fits-all approach, where all examinees answer the same set of questions, and a final score is calculated. Examples include traditional paper-and-pencil tests for humans and various gold-standard benchmarks for AI models.
However, as the testing scale increases and the complexity and diversity of agents grow, traditional assessment methods face challenges in efficiency and reliability. Computerized Adaptive Testing (CAT), originating from psychometrics, offers a personalized testing paradigm by identifying and presenting the most informative and valuable questions to each examinee [28, 163]. This method has been widely adopted in high-stakes testing scenarios for humans, such as the SAT, GRE, and GMAT [48, 107]. Recently, CAT has also been increasingly used to assess AI’s capabilities, such as textual entailment recognition, chatbots, machine translation, and general-purpose AI systems [124, 86, 141, 170]. CAT approach has been proved to require fewer questions to achieve the same level of assessment accuracy for both humans and AI systems [190, 81]. Essentially, CAT aims to address a critical question about accuracy and efficiency: How to accurately estimate an examinee’s true proficiency while minimizing the number of questions provided?
@@ 1 Introduction
It is a dynamic and interactive process between an examinee (human or AI model) and a testing system. The testing system includes four main components that take turns: At each test step, the Measurement Model, as the user model, first uses the examinee’s previous responses to estimate their current proficiency, based on cognitive science or psychometrics [1]. Then, the Selection Algorithm picks the next question from the Question Bank according to certain criteria [106, 23, 16]. Most traditional criteria are statistical informativeness metrics, e.g., selecting the question whose difficulty matches the examinee’s current proficiency estimate, meaning the examinee has roughly a 50% chance of getting it right. The above process repeats until a predefined stopping rule is met. Throughout the assessment, Test Control governs various factors such as exposure balance, fairness, and robustness of the testing. At the conclusion of CAT, the final proficiency estimate—or diagnostic report—serves as the outcome of the assessment.
@@
CAT represents a complex fusion of machine intelligence and assessment techniques. It needs to manage large question banks, adapt to varying examinee proficiencies, and real-time decision-making. Moreover, practical CAT also involves ensuring reliability, fairness, search efficiency, etc. These challenges make CAT a multifaceted decision-making problem. With the rise of large-scale and diverse online testing platforms, these challenges have become even more significant. Machine learning (ML), particularly deep learning, offers promising solutions to enhance both the efficiency and accuracy of testing. Previous CAT surveys [25, 163, 31, 118] have primarily focused on statistical and psychometric perspectives, concentrating mainly on human assessments. Given CAT’s interdisciplinary nature, this paper seeks to explore and review methodologies from a machine-learning perspective. It is more accessible to a broader readership and provides insights into building strong testing systems for both humans and artificial intelligence.
@@
In the realm of ML, CAT can be conceptualized as a parameter estimation problem with a focus on data efficiency [115, 191]: The objective is to determine the values of latent parameters within a model (i.e., the examinee’s true proficiency) using the minimum amount of observed data (i.e., the fewest possible questions answered by the examinee). In recent years, there has been a growing interest in applying ML techniques to investigate the four components in CAT. For example, deep learning techniques diagnose examinee’s proficiency [167] and automate question bank construction [126]; data-driven approaches optimize selection algorithms by learning from large-scale response data [189, 97, 57]. Despite these efforts, a comprehensive survey that captures the breadth of CAT solutions from a machine-learning perspective is still lacking. Furthermore, the ongoing evolution of machine learning presents new aspects for testing. The contributions of this paper are as follows:
@@
  • To our knowledge, this represents the first attempt to comprehensively review CAT solutions through the lens of machine learning. By exploring the existing work in Measurement Model, Selection Algorithms, Question Bank Construction, and Test Control, the paper offers a unified framework and encompasses the entire life cycle of the CAT system.
  • We summarize existing works and draw conclusions on the success and failure attempts of machine learning. Furthermore, we identify key factors that are essential for building reliable and effective CAT systems for both human and AI model evaluation, including exposure control, fairness, robustness, and search efficiency. It offers a more comprehensive perspective.
@@
The paper is organized as follows. In Section 2, we introduce the background of CAT. Then in Section 3, we provide the formulation of CAT’s task. After that, Section 4, 5, and 6 respectively review the existing methods for the measurement model, selection algorithms, and question bank construction. Given the fact that the selection algorithm is the core component for achieving the adaptivity, this survey mainly focuses on its recent advancements in machine learning and deep learning. In Section 7, we summarize the key factors in the application of CAT. Section 8 discusses how to evaluate the CAT.
2 Evolution of CAT
The evolution of CAT is a fascinating journey through time, marked by significant milestones. Adaptive testing began with Alfred Binet’s intelligence test in 1905 [165]. The 1950s saw the advent of computers, transforming adaptive testing into CAT. Key advancements in the 1970s and 1980s, particularly the integration of the psychometric model, enhanced assessment accuracy [138, 135]. The 1990s internet boom made CAT widely accessible, leading to its use in major tests like the GRE, GMAT, and SAT. These tests, though evolved, still rely on adaptive principles. Various statistical methods optimize the testing experience, and CAT became a major focus in human measurement, covering education [162, 107, 166], healthcare [60, 58, 59], sociology [117], and sports [4, 182].
Recently, researchers have increasingly explored applying CAT to AI model evaluation. Existing benchmarks often contain redundant, low-quality, contaminated, or even erroneous questions, affecting the efficiency and reliability of AI assessments [190]. By leveraging adaptive testing, researchers can analyze the characteristics of benchmark questions to customize assessments for each AI system and estimate the latent traits behind each model’s responses, rather than merely calculating accuracy. Guided by CAT and psychometrics, various efficient methods have emerged in various aspects of AI evaluation, including performance estimation [86, 125], question selection [63, 133, 190], and understanding experimental results [110, 111]. These methods aim to identify informative and valuable subsets from large-scale datasets to improve the reliability of AI system evaluations.
@@ 2 Evolution of CAT
Current CAT research spans a wide range of topics, including the development of question banks, question selection, proficiency estimation, and various issues related to test security and reliability. They are critical to ensuring that CAT remains a reliable, valid, and fair assessment. Machine Learning is revolutionizing CAT by enabling sophisticated analysis of large datasets, detailed behavior modeling, and flexible adaptation to diverse testing environments [189]. Despite the ML in CAT is still in its early stages, its potential is evident. Machine learning offers new solutions to improve how we define, analyze, and apply CAT [185]. This survey aims to provide an overview and understanding of traditional statistical-based and recent ML-based CAT.
@@
Figure 1: The workflow of CAT: At step $t$, the selection algorithm adaptively selects next question $q_{t+1}$ based on examinee’s current proficiency ${\theta}^{t}$ estimated by measurement models.
3 Overview
An important assumption [25] of CAT is that examinee’s true proficiency level $\theta_{0}\in\mathbb{R}^{d}$ is constant throughout the test. Here, $d$ represents the proficiency’s dimension; for example, $\theta$ may correspond to a unidimensional overall ability level ($d=1$) or a multidimensional vector representing mastery levels across $d$ distinct knowledge concepts. The primary goal of CAT is to accurately and efficiently estimate examinees’ true proficiency levels by having them answer questions. Thus, CAT systems are designed to achieve two key objectives: (1) to use the responses to estimate an examinee’s proficiency $\theta$ such that it closely approximates the true proficiency $\theta_{0}$ by the end of the test, and (2) to select the most valuable and fitting questions for each examinee, thereby reducing test length.
3.1 Task Formalization
To achieve the aforementioned objectives, CAT operates as an iterative and interactive process: As illustrated in Figure 1, at test step $t\in[1,2,...,T]$ in CAT, examinee’s current proficiency estimate $\hat{\theta}^{t}$ is estimated using previous $t$ responses; then leverage $\hat{\theta}^{t}$ to retrieve the next question $q_{t+1}$ from question bank $\mathcal{Q}$ to ask examinee, and receive the next response label $y_{t+1}$. These interactions form a response sequence $\{(q_{1},y_{1}),(q_{2},y_{2}),...,(q_{T},y_{T})\}$, where $y_{t}=1$ if the response to $q_{t}$ is correct and 0 otherwise. To achieve the goals of CAT, each test step involves two critical processes:
@@ 3.1 Task Formalization
(1) Proficiency Estimation. The Measurement Model, denoted by $f(\cdot)$, acts as a user model, predicting the probability of a correct response by an examinee with proficiency $\theta$, which is denoted as $f(q,\theta)=P(y=1|q,\theta)$. The implementation of measurement model often draws upon cognitive science [40] or psychometrics[1]. To accurately estimate examinee’s proficiency at each step, various estimation methods can be used, e.g., Maximum Likelihood Estimation (MLE) or Bayesian Estimation. In applications, the binary cross-entropy loss is frequently utilized: at step $t$, given previous $t$ responses $\mathcal{D}_{1:t}=\{(q_{1},y_{1}),(q_{2},y_{2}),...,(q_{t},y_{t})\}$, the corresponding empirical loss is:
${L}(\theta)$ $=\sum_{(q,y)\in\mathcal{D}_{1:t}}\ell(y,f(q,\theta))$ (1)
$=-\sum_{(q,y)\in\mathcal{D}_{1:t}}{y\log f(q,\theta)+(1-y)\log(1-f(q,\theta))},$
@@
thus the current estimate of proficiency, $\hat{\theta}^{t}$, is obtained by minimizing the loss function $L(\theta)$: $\hat{\theta}^{t}=\mathop{\arg\min}_{\theta}{{L}(\theta)}$.
(2) Question Selection. The heart of CAT is an algorithm that picks the next question $q_{t+1}$ from the question bank $\mathcal{Q}$, using examinee’s current proficiency estimate $\hat{\theta}^{t}$ as a guide:
$q_{t+1}=\mathop{\arg\max}_{q\in\mathcal{Q}}\mathcal{V}_{q}(\hat{\theta}^{t}),$ (2)
@@
where $\mathcal{V}_{q}(\hat{\theta}^{t})$ is the value of question $q$. For instance, $\mathcal{V}$ might be a measure of how much information the question will provide about the examinee’s proficiency, or it could be the output of a policy $\pi$ specifically designed to determine question selection.
After receiving new response label $y_{t+1}$, measurement model updates and estimates proficiency $\hat{\theta}^{t+1}$. The above process will be repeated for $T$ times, ensuring the final step estimate $\hat{\theta}^{T}$ close to the true $\theta_{0}$, i.e.,
Definition 1 (Definition of CAT)
@@
$\min_{|S|=T}\|{\hat{\theta}^{T}}-\theta_{0}\|.$ (3)
However, solving this optimization problem directly is impractical, as the true proficiency $\theta_{0}$ is not observable, and even the examinees may not know their exact proficiency level. Consequently, existing methods are all approximations of this target. For example, traditional statistical selection methods [106, 23] utilize the asymptotic statistical properties of the MLE to reduce estimation uncertainty, e.g., selecting questions whose difficulty closely match the examinee’s current estimated proficiency $\hat{\theta}^{t}$. More recent Subset Selection approaches [191] try to identify a theoretical approximation of $\theta_{0}$ to serve as a new objective for optimization. For further details, refer to Section 5.
Meanwhile, as a practical system, considerations extend beyond proficiency estimation objective (Definition 1). Factors such as question exposure control, robustness, fairness, and search efficiency must be addressed as well. An exhaustive discussion of these factors is presented in Section 7.
Evaluation Methods: To validate the accuracy of the estimated proficiency, two primary approaches are employed: 1) Performance prediction: using the examinee’s estimated values within the measurement model to predict the correctness label $y$ of the responses on examinee’s reserved response data, often measured by cross-entropy; 2) Proficiency estimation: using simulation to generate true proficiency values $\theta_{0}$, simulating the examinee’s responses to each question. Then the Mean Squared Error (MSE) between the estimates and the simulated true values can be calculated. The details can be found in Section 8.
@@
Figure 2: Summary of representative Computerized Adaptive Testing methods in machine learning perspective.
3.2 Categorization